OpenStudy (mony01):

If f(x) is continuous on (0,1), then it reaches both its absolute maximum and minimum somewhere inside (0,1). True or False

OpenStudy (ddcamp):

What does being continuous mean? Could you have a function that is continuous on (0,1) that looks like this: 

OpenStudy (mony01):

I know that f(x) is continuous on closed interval [0,1], but im not sure if its continuous on (0,1). Im not sure if the question is trying to trick me or not.

OpenStudy (ddcamp):

Are you given a function? Or is this just a general question?

OpenStudy (mony01):

just a true or false question

OpenStudy (anonymous):

No, the extreme value theorem is only valid for compact sets. Since \((0, 1)\) is not compact the extreme value theorem does not hold.

OpenStudy (anonymous):

We can provide a counterexample to show this. Consider \(f(x) = \frac{1}{x}\) on the interval \((0, 1)\). It is continuous on that interval but it certainly does not attain a maximum.

OpenStudy (anonymous):

For more information see http://en.wikipedia.org/wiki/Extreme_value_theorem

OpenStudy (mony01):

i think we used the mean value theorem

OpenStudy (anonymous):

Yes... but this question does not involve the mean value theorem. Read the statement of the extreme value theorem carefully.

OpenStudy (anonymous):

If we are working with functions of the form \(f : \mathbb{R} \to \mathbb{R}\) then the extreme value theorem states that \(f\) is continuous and consider \([a, b]\) a closed and bounded interval then \(f\) attains a maximum and minimum value on that interval.

OpenStudy (mony01):

so can it be continuous on both closed and open interval?

OpenStudy (anonymous):

No, that's the reason it doesn't work in the example you provided. Because its an open interval.

OpenStudy (anonymous):

The extreme value theorem only applies to closed and bounded intervals.

OpenStudy (mony01):

oh okay I see now thanks.